# Tapas are free! Training‑Free Adaptation of Programmatic Agents via LLM‑Guided Program Synthesis in Dynamic Environments

### v1

@@ 5.3 Ablation Study
Table 4: Multi-Round Ablation on Formation Adaptation
@@ A.1 Autonomous cyber defense
We implement three DDoS attack patterns in the NS3 simulator to evaluate the effectiveness of action adaptation. DDoS attacks overwhelm target servers by coordinating multiple compromised nodes to generate malicious traffic that exhausts network resources or computational capacity (Zargar, Joshi, and Tipper 2013). Traditional automated network defense systems face significant challenges when confronting these attacks due to scalability constraints, parameter uncertainty, and evolving attack patterns. Static defense mechanisms struggle with dynamic scaling requirements as attack volumes fluctuate, face difficulties in determining optimal detection thresholds without prior knowledge of attack characteristics, and exhibit brittleness when encountering novel attack vectors that deviate from predefined setting.
@@
Each environment scenario E1-E5 operates with a 5-second simulation duration serving as the performance evaluation window. This temporal design balances real-time attack detection requirements with defense strategy adaptation latency, ensuring sufficient data samples for statistical significance analysis while maintaining practical deployment time constraints. TAPA’s meta-agent employs a TiT neuro-symbolic architecture (Mao et al. 2022), trained in baseline environment E1 to learn the mapping strategy from network state space to logical primitives. The training process establishes semantic foundations for symbolic defensive actions, providing a stable decision-making framework for subsequent action space adaptation. We incorporate expert knowledge such as MITRE ATT&CK knowledge base (Strom et al. 2018) to guide Claude Sonnet 4, providing domain-specific guidance principles for symbolic program synthesis. When performance degradation triggers are detected within the 20-second evaluation window (network uptime $50\%$ or active flow count exceeding three times the baseline), we transmit current environmental context, performance variations, and historical adaptation experiences from previous code modifications to Claude Sonnet 4 for symbolic program synthesis execution.
@@ A.2 Swarm formation control
TAPA’s meta-agent operates over two logical primitives based on (Hu et al. 2025a): L1 (Formation Control) for adjusting coordination parameters such as cohesion distance and behavioral weights, and L2 (Defend) for implementing defensive programs including outlier detection and adversarial resilience mechanisms. The meta-agent autonomously selects appropriate logical primitives based on formation quality metrics and environmental context indicators. When formation consensus fails within 150 simulation frames, GPT-4o synthesizes adaptive control programs incorporating domain expertise in distributed control theory and multi-agent coordination principles. This enables real-time adaptation to both environmental changes and adversarial interference while maintaining interpretability and formal guarantees essential for safety-critical swarm applications.

### v2

@@ 5.3 Ablation Study
Table 4: Multi-Round Ablation on Formation Adaptation
@@ 6 Conclusion
Acknowledgments
This work is partially funded by the European Union (under grant agreement ID 101212818). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HADEA). Neither the European Union nor the granting authority can be held responsible for them. This work is partially supported by Innovate UK through AI-PASSPORT under Grant 10126404. Yi’s contribution is partially supported through the Royal Society international exchanges programme and in part by the Engineering and Physical Sciences Research Council, through funding from RAi UK [EP/Y009800/1].
@@ A.1 Autonomous cyber defense
We implement three DDoS attack patterns in the NS3 simulator to evaluate the effectiveness of action adaptation. DDoS attacks overwhelm target servers by coordinating multiple compromised nodes to generate malicious traffic that exhausts network resources or computational capacity (zargar2013survey). Traditional automated network defense systems face significant challenges when confronting these attacks due to scalability constraints, parameter uncertainty, and evolving attack patterns. Static defense mechanisms struggle with dynamic scaling requirements as attack volumes fluctuate, face difficulties in determining optimal detection thresholds without prior knowledge of attack characteristics, and exhibit brittleness when encountering novel attack vectors that deviate from predefined setting.
@@
Each environment scenario E1-E5 operates with a 5-second simulation duration serving as the performance evaluation window. This temporal design balances real-time attack detection requirements with defense strategy adaptation latency, ensuring sufficient data samples for statistical significance analysis while maintaining practical deployment time constraints. TAPA’s meta-agent employs a TiT neuro-symbolic architecture (mao2022transformer), trained in baseline environment E1 to learn the mapping strategy from network state space to logical primitives. The training process establishes semantic foundations for symbolic defensive actions, providing a stable decision-making framework for subsequent action space adaptation. We incorporate expert knowledge such as MITRE ATT&CK knowledge base (strom2018mitre) to guide Claude Sonnet 4, providing domain-specific guidance principles for symbolic program synthesis. When performance degradation triggers are detected within the 20-second evaluation window (network uptime $50\%$ or active flow count exceeding three times the baseline), we transmit current environmental context, performance variations, and historical adaptation experiences from previous code modifications to Claude Sonnet 4 for symbolic program synthesis execution.
@@ A.2 Swarm formation control
TAPA’s meta-agent operates over two logical primitives based on (hu2025enhancingrobustnessllmdrivenmultiagent): L1 (Formation Control) for adjusting coordination parameters such as cohesion distance and behavioral weights, and L2 (Defend) for implementing defensive programs including outlier detection and adversarial resilience mechanisms. The meta-agent autonomously selects appropriate logical primitives based on formation quality metrics and environmental context indicators. When formation consensus fails within 150 simulation frames, GPT-4o synthesizes adaptive control programs incorporating domain expertise in distributed control theory and multi-agent coordination principles. This enables real-time adaptation to both environmental changes and adversarial interference while maintaining interpretability and formal guarantees essential for safety-critical swarm applications.